OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) = 3x2 + 12x + 16 g(x) = 2 sin(2x - π) + 4 Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value.

OpenStudy (dls):

\[\LARGE f(x)= 3x^2 + 12x + 16 \] To find the minimum value of f(x),We use calculus and differentiate it n equate to 0. Let it be equal to "y".\[\Large \frac{dy}{dx}=6x+12=0 =>x=-2\]

OpenStudy (dls):

Similarly for g(x)\[\LARGE g(x) = 2 \sin(2x - π) + 4\] \[\Large \frac{dy}{dx}=2 \cos(2x-\pi) \times 2 = 0 =>2x-\pi = \frac{\pi}{2}\] \[\LARGE x = \frac{3 \pi}{4}\]

OpenStudy (dls):

Now you need to put both the values of x in both the equations respectively to find the minimum value and compare which is lesser as asked in the question.

OpenStudy (dls):

You can find the 2nd derivative of both the functions to insure that \[\Huge \frac{d^2y}{dx^2}>0\] hence minimima

OpenStudy (anonymous):

the function f(x) is an equation of parabola and it has its minimum value for x=-2 and f(x)=4 while the other function g(x) has its minimum value of 2 ... as sine has a smallest value of -1

OpenStudy (anonymous):

What are the values of x

OpenStudy (dls):

-2 and 3pi/4

OpenStudy (anonymous):

How did you get them?

OpenStudy (anonymous):

Sorry I just don't understand these type of problems